Seismic Analysis

1) Define interstorey drift ratio (IDR). How do we find IDR and how do we find the maximum IDR. Whats the difference between IDR and maximum IDR. What is the significance of IDR and maximum IDR. If there is a standard for these values please do mention the clause for me.

Hi there, please clarify the above things for me, the example i have is a 4 storey building so will be helpful if you can explain to me using a example as well. Thanks alot :)
Answer #1

Ans - When the displacements of the roof and the floor of any given storey when there is sway of the building due to some seismic activity, then the difference between these displacements normalized by the storey height is called Interstorey drift (IDR)

IDR depends on the direction of sway

= Drift at a given storey/ height from that storey to the storey below

The IDR max is always maximum at ground storey level while normal IDR in general is calculated at any other storey.So exclusively IDR will be maximum at 1st level.

But IDR at ground is called GSDR.

Suppose if drift is X and heights be H

1st level drift = X1

2nd level drift = X2 - X1

Drift ratio @ 1 = X1 / H1

@ 2 = (X2 - X1) / H2